OpenStudy (rayford):

Choose the value of the coefficient determinant (D) in the following system. 2x – 5y = –2 x – 3y = 0

OpenStudy (anonymous):

The coefficient matrix is made up of the coefficients of x and y: \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right] = \left[\begin{matrix}2 & -5 \\ 1 & -3\end{matrix}\right]\] The determinant of that is found by multiplying the diagonal terms and subtracting them in the form of ad-bc: \[\left| \left[\begin{matrix}2 & -5 \\ 1 & -3\end{matrix}\right] \right| = 2(-3)-(-5)(1)=-6+5=-1\]

OpenStudy (rayford):

Choose the value of the x determinant (Dx) in the following system. 3x – 4y = –7 x + y = 7

OpenStudy (rayford):

their is an x?? instead of just an an x

OpenStudy (rayford):

d i mean

OpenStudy (anonymous):

Hmm I am not sure what is meant by "x determinant." I don't believe I've heard that term before. However here is a link to a question similar to yours, perhaps the answer there can help you: http://answers.yahoo.com/question/index?qid=20101219091339AAmbZLk

OpenStudy (anonymous):

What are the choices of values you are supposed to choose from?

OpenStudy (rayford):

21 28 -14 -35

OpenStudy (anonymous):

Ok if you use the method the person used in that link I posted, you will get a number that is included in those choices you listed.

OpenStudy (rayford):

its 35! :[)?

OpenStudy (anonymous):

Well that is not what I got. Show me what you did.

OpenStudy (rayford):

[3.4] [1.1] 4-3=1 -7.4 7.1 28+7=35??

OpenStudy (anonymous):

Your answer is a little confusing. Why are you doing 4-3=1? Note: your "D" coefficient matrix needs to have a -4, not 4.

OpenStudy (rayford):

so the answer is -35?

OpenStudy (anonymous):

No, how would you get -35? Plug in the -4 in and see what you get when you do the math. \[\left[\begin{matrix}-7 & -4 \\ 7 & 1\end{matrix}\right] = ?\]

OpenStudy (anonymous):

(-7)(1)-(-4)(7)=-7+28 = ...?

OpenStudy (rayford):

ohhhhhh....thats how i got 35..i got the negatives mixed up..

OpenStudy (rayford):

its 21 ok..

OpenStudy (anonymous):

Yep, just had to have a -4 instead of a 4. If the method that person in the link posted is correct, then your answer should be correct.

OpenStudy (rayford):

may you do one more?

OpenStudy (rayford):

Choose the value of the y determinant (Dy) in the following system. 2x – 4y = 24 2x – 3y = 15 –132 –12 –18 78

OpenStudy (anonymous):

Well now that you have done one, you should be able to do another one, using that same method. I assume now that it's y instead of x though, you use the x column (2,2) instead of the y column like in the previous problem.

OpenStudy (rayford):

well can you atleast check my answer :3?

OpenStudy (anonymous):

Sure.

OpenStudy (rayford):

-12?

OpenStudy (anonymous):

Well, you get -12 if you use the y coefficients (-4,-3) like in the previous problem, which asked you for the x determinant, so I assume that -12 is your x determinant. If you want the y determinant, I assume you have to use the x coefficients then (2,2)... but that gives me an answer of 18, not -18 like in your choices. So I am uncertain here -- hopefully someone else can pick it up.

OpenStudy (rayford):

it was 18..... thank you and goodnight

OpenStudy (anonymous):

Oh okay. You are welcome. Good night :)